[figure: 9.7 cm]

NUCLEAR MAGNETIC RESONANCE OF ^{57}Fe IN BISMUTH-YTTRIUM GARNETS

H. Štěpánková^{a}, T. Okuda^{b}, E. Gamaliy^{a}, A. Snezhko^{a}, N. Adachi^{b}, T. Kaneuji^{b} aFaculty of Mathematics and Physics, Charles University 180 00 Prague 8, V Holešovičkách 2, Czech Republic bNagoya Institute of Technology Gokiso-cho, Showa-ku, Nagoya,466-8555 Japan

We report ^{57}Fe NMR spectra of a bismuth-yttrium iron garnet system Bi_{x}Y_{3−x}Fe_{5}O_{12}. The spectra were measured using spin echo technique at 4.2 K in zero external magnetic field. We were able to detect the ^{57}Fe NMR signal of a single crystal film of bismuth iron garnet Bi_{3}Fe_{5}O_{12} prepared by ion beam sputtering on Nd_{3}Ga_{5}O_{12} substrate. The spectrum is in agreement with iron garnet structure having magnetization along [111] direction: one d-line at 66.46 MHz assigned to tetrahedral iron sites a-lines at 74.00 and 75.05 MHz with the intensity ratio 1:3 assigned to octahedral iron sites. The resonant frequencies are slightly shifted in comparison with yttrium iron garnet Y_{3}Fe_{5}O_{12}. Spectra of partially substituted samples (x=0.16 and 2.75) exhibit satellite structure or broadening corresponding to the distribution of Y and Bi in dodecahedral sites.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : H. Štěpánková

Address for correspondence : Faculty of Mathematics and Physics Charles University 180 00 Prague 8 V Holešovičkách 2 Czech Republic

Email address : stepanko@mbox.troja.mff.cuni.cz